### v2

# thanks: Corresponding authorthanks: Corresponding author

Abstract
[…]
I INTRODUCTION
The simplest $\Lambda$ cold dark matter ($\Lambda\mathrm{CDM}$) model has been remarkably consistent with most astronomical observations across different epochs and scales of the Universe Riess et al. (1998); Perlmutter et al. (1999); Brout et al. (2022); Asgari et al. (2021); Troxel et al. (2018); Aiola et al. (2020). However, the persistent discrepancy in the value of the Hubble constant $H_{0}$ inferred from the early and late Universe, known as the Hubble tension, is challenging the $\Lambda\mathrm{CDM}$ model Di Valentino et al. (2021a); Perivolaropoulos and Skara (2022); Shah et al. (2021); Verde et al. (2019). Among various discrepancies, the most significant is the over $5\sigma$ divergence between the Planck 2018 estimate of $H_{0}=67.4\pm 0.5\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$, indirectly inferred from the cosmic microwave background (CMB) data assuming the $\Lambda\mathrm{CDM}$ model Aghanim et al. (2020), and the SH0ES result of $H_{0}=73.04\pm 1.04\,\mathrm{km\ s^{-1}\ Mpc^{-1}}$ directly measured from the type Ia supernovae (SNe) calibrated using the Cepheid variables Riess et al. (2022). The tension cannot be merely explained by a statistical fluctuation, even after the reanalysis of systematic uncertainties Akrami et al. (2020); Zhang et al. (2017a). Therefore, the irreconcilable tension may indicate underlying systematics or new physics beyond $\Lambda\mathrm{CDM}$ Di Valentino et al. (2021b); Bernal et al. (2016); Abdalla et al. (2022); Di Valentino et al. (2025).
However, no consensus has yet been reached on resolving the tension based on current data. Given this discrepancy between the early and late Universe, proposed scenarios can broadly be classified into early-time or late-time solutions.
In early-time solutions, new physics is introduced to reduce the sound horizon, $r_{\mathrm{d}}$, thereby increasing $H_{0}$ while preserving the baryon acoustic oscillation (BAO) measurement of $r_{\mathrm{d}}H_{0}$. One realization involves modifying expansion history by introducing additional energy components around the epoch of matter-radiation equality, such as early dark energy Poulin et al. (2019); Karwal and Kamionkowski (2016); Mörtsell and Dhawan (2018); Kamionkowski et al. (2014); Berghaus and Karwal (2020) or extra neutrino species Zhang et al. (2015a); Zhao et al. (2017a, b); Feng et al. (2020); Kreisch et al. (2020); Wyman et al. (2014); Sakstein and Trodden (2020); Zhang et al. (2014a). Alternatively, early-time solutions may also involve modifying the recombination history Sekiguchi and Takahashi (2021); Chiang and Slosar (2018); Mirpoorian et al. (2025) or introducing primordial magnetic fields Jedamzik and Pogosian (2020); Jedamzik et al. (2025); Jedamzik and Abel (2013). However, these early-time solutions are in tension with other cosmological observables, such as the fluctuation amplitude $\sigma_{8}$ [^1] However, using $\sigma_{8}$ may bias the comparison between models with high $H_{0}$ and $\Lambda\mathrm{CDM}$, since the reference scale $R=8h^{-1}~\mathrm{Mpc}$ depends on $h$, which changes the interpretation of $\sigma_{8}$ Forconi et al. (2025); Sanchez (2020). and the integrated Sachs-Wolfe effect, as discussed in the no-go arguments of Refs. Hill et al. (2020); Krishnan et al. (2020); Jedamzik et al. (2021); Vagnozzi (2023); Dutta et al. (2019). Late-time solutions modify the Hubble parameter $H(z)$ in the late universe, so that it can accommodate the SH0ES $H_{0}$ result locally while leaving the CMB anisotropy spectrum unchanged, such as the deformation of the dark-energy equation of state Wang et al. (2016); Agrawal et al. (2023); Li and Shafieloo (2019); Li (2004); Shapiro and Sola (2002); Zhang et al. (2012); Li et al. (2013); Zhang et al. (2015b); Feng and Zhang (2016), modified gravity Aluri et al. (2023); Capozziello and De Laurentis (2011); De Felice and Tsujikawa (2010), and parameter transitions in the late Universe Mortonson et al. (2009); Kazantzidis and Perivolaropoulos (2019). Analogous to early-time solutions, various no-go arguments have also been raised against the late-time scenarios. The requirement of consistency between low redshift cosmological distance measurements from BAO and SNe imposes strong constraints on the late-time solutions Benevento et al. (2020); Yang et al. (2021). Moreover, these solutions cannot reconcile the $H_{0}$ tension and $S_{8}$ tension [^2] The $S_{8}$ tension is alleviated as indicated by the updated Kilo-Degree Survey cosmic shear constraints with improved redshift distribution estimation and calibration, which found consistency ($0.73\sigma$) with Planck Wright et al. (2025). simultaneously, as Refs. Guo et al. (2019); Gao et al. (2021); Heisenberg et al. (2022); Dutta et al. (2019); Alestas and Perivolaropoulos (2021) pointed out.
In light of the plethora of proposed cosmological models and corresponding no-go arguments, it is compelling to explore new and independent measurements of $H_{0}$. For instance, techniques such as strong gravitational lensing time delay Birrer et al. (2019); Shajib et al. (2020) and gravitational wave standard sirens Guidorzi et al. (2017); Song et al. (2025); Jin et al. (2025) provide an independent perspective on this tension. We can also employ other independent calibrators to replace Cepheid variables in the distance ladder, such as megamasers Pesce et al. (2020), the tip of the red giant branch Freedman (2021), surface brightness fluctuations Blakeslee et al. (2021), and Mira variables Huang et al. (2019).
The inverse distance ladder (IDL) calibrates the SNe via anchored BAO instead of Cepheid variables, providing a model-insensitive method to infer $H_{0}$ Aubourg et al. (2015); Cuesta et al. (2015); Verde et al. (2017); Lemos et al. (2019); Yu et al. (2018); Gómez-Valent and Amendola (2018); Haridasu et al. (2018); Gómez-Valent et al. (2024). In the traditional IDL, the BAO data are anchored using an $r_{\mathrm{d}}$ prior derived from the CMB or big bang nucleosynthesis. Then the data are used to calibrate SNe to give an absolute distance. Finally, the distance is extrapolated to $z=0$ to infer $H_{0}$ under a fiducial $H(z)$ model Percival et al. (2010); Heavens et al. (2014); Abbott et al. (2018); Alam et al. (2021); Camilleri et al. (2025); Adame et al. (2025); Macaulay et al. (2019).
It should be noted that the $r_{\mathrm{d}}$ prior introduces dependence on early-Universe physics Aylor et al. (2019). The effect of this $r_{\mathrm{d}}$ prior has been systematically evaluated in Refs. Barua and Desai (2025); Luongo and Muccino (2024). To mitigate such dependence, the $r_{\mathrm{d}}$ prior can be replaced with other independent measurements. The cosmic chronometer (CC) measures $H(z)$ directly, rather than through integrals of $H(z)$, such as those involved in lensing distances from strong lensing systems Wojtak and Agnello (2019); Taubenberger et al. (2019); Wang et al. (2022); Gong et al. (2024), thus offering higher sensitivity to cosmological parameters Moresco (2024); Vagnozzi et al. (2022); Jimenez et al. (2019). CC data have been widely applied as an effective replacement for $r_{\mathrm{d}}$ in constructing IDL Guo et al. (2025); Luongo and Muccino (2024); Favale et al. (2023); Cai et al. (2022a).
Furthermore, adopting a fiducial $H(z)$ derived from a specific cosmological model in IDL will inevitably introduce model dependence. To avoid this, the traditional approach is to parametrize $H(z)$ via a Taylor expansion in terms of $z$ or $y\equiv 1-a=z/(1+z)$, known as the cosmography method Camilleri et al. (2025); Pourojaghi et al. (2025); Luongo and Muccino (2024); Lemos et al. (2019); Cattoen and Visser (2007); Zhang et al. (2017b); Chiba and Nakamura (1998); Capozziello et al. (2011); Aviles et al. (2012). Alternatively, purely data-driven techniques such as the Gaussian process (GP) have been employed to reconstruct $H(z)$ Guo et al. (2025); Jiang et al. (2024); He et al. (2022); Mukherjee and Sen (2024); Yang et al. (2023). However, both approaches possess intrinsic limitations. The Taylor expansion is problematic at $z\sim 1$ Banerjee et al. (2021); Cai et al. (2022a, b); Zhang et al. (2017b). GP methods are sensitive to the choice of kernel functions and tend to yield smaller errors at the cost of strongly correlated results Ó Colgáin and Sheikh-Jabbari (2021); Kjerrgren and Mortsell (2022). To address these limitations, we adopt a global parametrization based on the cosmic age (PAge) Huang (2020); Luo et al. (2020) and its extension, a more accurate parametrization based on cosmic age (MAPAge) Cai et al. (2022b); Huang et al. (2021a), to accurately reproduce late-time cosmological behaviors.
The PAge model can be interpreted as an effective description of homogeneous late-time scenarios. The studies Huang et al. (2021b, 2022); Li et al. (2022) have employed the PAge model to analyze gamma-ray bursts, redshift-space distortions, and quasars. References Cai et al. (2022a, b); Huang et al. (2025a) have derived no-go theorems against late-time solutions through the comparison between the PAge (denoted as late-time scenarios) and $\Lambda\mathrm{CDM}$ models.
Recently, the Dark Energy Spectroscopic Instrument (DESI) Collaboration reported a $3.9\sigma$ hint of dynamical dark energy in the first data release (DR1), which was further increased to $4.2\sigma$ in the second data release (DR2) Adame et al. (2025); Abdul Karim et al. (2025). This finding has sparked intense discussions about possible new physics and systematic uncertainties Malekjani et al. (2025); Colgáin and Sheikh-Jabbari (2024); Jia et al. (2025); Wang et al. (2024a); Huang et al. (2025b); Cortês and Liddle (2024); Efstathiou (2025); Abreu and Turner (2025); Pang et al. (2025a); Fikri et al. (2025); Feng et al. (2025a); Du et al. (2025a); Li et al. (2025a, 2024, b); Du et al. (2025b); Pang et al. (2025b); Wang et al. (2025a); Wang and Piao (2024); Lodha et al. (2025); Pan et al. (2025); You et al. (2025); Pan and Ye (2025); Silva et al. (2025); Qiang et al. (2025); Cai et al. (2025); Cline and Muralidharan (2025); Rodrigues et al. (2025); Li and Wang (2025); Bhattacharjee et al. (2025); Gialamas et al. (2025a); Wang and Piao (2025); Cline and Muralidharan (2025); Li et al. (2025c); Li and Zhang (2025); Li et al. (2025d, e); González-Fuentes and Gómez-Valent (2025); Gómez-Valent and Solà Peracaula (2025); Wang et al. (2025b); Du et al. (2025c); Feng et al. (2025b). This behavior, which deviates from $\Lambda\mathrm{CDM}$ can be effectively described by the PAge model. This motivates us to propose a PAge-improved IDL to yield fully model-independent cosmological constraints in this work. Our results will provide model-independent perspectives on the $H_{0}$ tension and the hint of dynamical dark energy.
The remainder of this paper is organized as follows. In Sec. II, we introduce the PAge model and cosmological datasets used in our analysis. We present the cosmological constraints and related discussion in Sec. III. The paper concludes with a summary in Sec. IV.
II METHODOLOGY
II.1 The PAge and MAPAge parametrization
Within the IDL, the PAge (MAPAge) parametrization is used to extrapolate $H(z)$ to $z=0$, thereby determining $H_{0}$ in a cosmological-model-independent manner. The PAge (MAPAge) rests on two assumptions:
  1. (i)
    At $z\gg 1$, the Universe is matter dominated, and the brief radiation-dominated era is neglected.
  2. (ii)
    In PAge, the product of the Hubble parameter $H$ and the cosmic time $t$, can be approximated by a quadratic function of $t$. In MAPAge, this product can be approximated by a cubic function of $t$.
These assumptions are independent of any specific cosmological model, such as dark energy models and modified gravity theories Huang (2020); Cai et al. (2022a, b); Huang et al. (2025a); Wang et al. (2024b). Rather than parametrizing the dark energy equation of state, PAge directly parametrizes the cosmic expansion history, providing a more straightforward probe of that history than equation of state based approaches such as $w_{0}w_{a}\mathrm{CDM}$. Compared with traditional parametrizations, such as expansions in $z$ or $y$ Capozziello et al. (2011); Zhang et al. (2017b); Visser (2004), PAge (MAPAge) can parametrize $H(z)$ over a wider redshift range with higher accuracy. As shown in Fig. 1, PAge (MAPAge) can approximate the target $w_{0}w_{a}\mathrm{CDM}$ model up to $z\sim 10^{3}$, whereas $z$ and $y$ expansion develop deviation at $z>1$ Hu and Wang (2022); Cai et al. (2022a). Moreover, for target models where $H(z)$ evolves rapidly, a Taylor expansion in terms of $z$ or $y$ requires high-order terms to accurately capture the behavior. However, since $t$ is defined through the integral of $H(z)$, it evolves more smoothly. As a result, even pronounced changes in $H(z)$ can typically be captured by a low-order expansion in $t$. Thus, the PAge (MAPAge) model serves as a global, model-independent parametrization that can be effectively adopted in IDL.
In PAge, we first rewrite $H(t)t$ as
$H(t)t=H(t)(t-t_{0})+H(t)t_{0},$ (1)
where $t_{0}$ denotes the current cosmic age. Then we perform a Taylor expansion of $H(t)$ around $t_{0}$,
$H(t)=H_{0}+\left.\dot{H}\,\right|_{t=t_{0}}(t-t_{0})+\ldots.$ (2)
Substituting Eq. (2) into Eq. (1) yields
$H(t)t=$ $H_{0}t_{0}+H_{0}(t-t_{0})(1-H_{0}t_{0}-H_{0}t_{0}q_{0})$
$-H^{2}_{0}(t-t_{0})^{2}(1+q_{0}-\frac{H_{0}t_{0}}{2}(2+3q_{0}+j_{0}))+\dots.$ (3)
The deceleration and jerk parameters, $q_{0}$ and $j_{0}$, arise from the Taylor expansion of $H(t)$ in Eq. (2). As required by assumption (ii), we keep terms up to quadratic order in Eq. (3). To meet the assumption (i), $H(t)t=2/3$ in the limit $t\to 0$, we introduce the following modification of our expansion,
$\lim\limits_{t\to 0}H(t)t=H_{0}t_{0}+H_{0}(0-t_{0})(1-H_{0}t_{0}-H_{0}t_{0}q_{0})-AH^{2}_{0}t^{2}_{0},$ (4)
where $A$ is an undetermined parameter. The limit $\lim\limits_{t\to 0}H(t)t=2/3$ is met only if $A=1+q_{0}-\frac{2}{3H_{0}^{2}t^{2}_{0}}$. Therefore, the final expansion becomes
$H(t)t=$ $H_{0}t_{0}+H_{0}(t-t_{0})(1-H_{0}t_{0}-H_{0}t_{0}q_{0})$
$-H^{2}_{0}(t-t_{0})^{2}(1+q_{0}-\frac{2}{3H_{0}^{2}t^{2}_{0}}).$ (5)
After setting
$p_{\mathrm{age}}=H_{0}t_{0},~\eta=1-\frac{3}{2}p^{2}_{\mathrm{age}}(1+q_{0}),$ (6)
the final expansion is written as
$\frac{H(t)}{H_{0}}=1+\frac{2}{3}(1-\eta\frac{H_{0}t}{p_{\mathrm{age}}})(\frac{1}{H_{0}t}-\frac{1}{p_{\mathrm{age}}}),$ (7)
where $\eta$ quantifies the extent to which the real universe deviates from the pure matter universe with $\eta=0$ in Eq. (7). To obtain $z(t)$, we can replace $H(t)$ in Eq. (7) with its definition $H(t)=-\mathrm{d}z/\mathrm{d}t/(1+z)$, and integrate both sides,
$1+z=\left(\frac{p_{\mathrm{age}}}{H_{0}t}\right)^{2/3}\mathrm{e}^{\frac{1}{3}\left(1-\frac{H_{0}t}{p_{\mathrm{age}}}\right)\left(3p_{\mathrm{age}}+\eta\frac{H_{0}t}{p_{\mathrm{age}}}-\eta-2\right)}.$ (8)
Again, in Eq. (8), $\eta$ quantify the deviation from the matter-dominated Universe $1+z=(t_{0}/{t})^{\frac{2}{3}}$.
Similarly, we can parametrize $H(t)t$ with cubic function of $t$, which is referred to as MAPAge in Ref. Huang et al. (2021a). It takes the following forms:
$\frac{H}{H_{0}}=1+\frac{2}{3}$ $\left[1-(\eta_{1}+\eta_{2})\frac{H_{0}t}{p_{\mathrm{age}}}+\eta_{2}\left(\frac{H_{0}t}{p_{\mathrm{age}}}\right)^{2}\right]$
$\times\left(\frac{1}{H_{0}t}-\frac{1}{p_{\mathrm{age}}}\right),$ (9)
where $\eta_{1}$ equals to $\eta$ in PAge. $\eta_{2}$ is introduced to further parametrize $H(t)t$ and is given by
$\eta_{2}=1-\frac{3}{4}p^{2}_{\mathrm{age}}(2+j_{0}+3q_{0}),$ (10)
Figure 1: Comparison of approximation accuracy for PAge (green dash-dotted), MAPAge (red dashed), Taylor expansion in $z$ (brown dotted), and $y$ (light beige), targeting the fiducial $w_{0}w_{a}\mathrm{CDM}$ (black) model. The lower-right inset figure shows normalized Hubble parameter $H(z)/H(z)_{w_{0}w_{a}\mathrm{CDM}}$ up to $z\sim 10^{3}$. We fix the parameters of the $w_{0}w_{a}\rm CDM$ model to DESI DR2 results: $\Omega_{\mathrm{m}}=0.32,~H_{0}=66.74~\mathrm{km~s^{-1}~Mpc^{-1}},~w_{0}=-0.75,~w_{a}=-0.86$ Abdul Karim et al. (2025).
The PAge model can map specific dark energy models into a two-parameter space $(p_{\mathrm{age}},~\eta)$. Thus, the deviation from the $\Lambda\mathrm{CDM}$ model is clearly illustrated in the $(p_{\mathrm{age}},~\eta)$ parameter space. For simplicity, the mapping procedure is carried out by equating the deceleration parameter $q_{0}$ between the PAge model and the specific cosmological model, followed by computing $p_{\mathrm{age}}=H_{0}t_{0}$ for the reference model. Subsequently, $\eta$ can be derived from Eq. (6). More details about model mapping are provided in Ref. Luo et al. (2020).
In the following, we explicitly demonstrate how to map the specific dark energy models discussed in this paper into $(p_{\mathrm{age}},~\eta)$, including $\Lambda\mathrm{CDM}$, $w\mathrm{CDM}$, and $w_{0}w_{a}\mathrm{CDM}$.
For $\Lambda\mathrm{CDM}$ with $H(a)=H_{0}\sqrt{\Omega_{\mathrm{m}}a^{-3}+(1-\Omega_{\mathrm{m}})}$ when mapped to the PAge model, the deceleration parameter $q_{0}$ is
$q_{0}=-1-\frac{\dot{H}}{H^{2}_{0}}=-1+\frac{3}{2}\Omega_{\mathrm{m}},$ (11)
and $p_{\mathrm{age}}$ reads
$p_{\mathrm{age}}=\int^{1}_{0}\frac{\mathrm{d}a}{a\sqrt{\Omega_{\mathrm{m}}a^{-3}+(1-\Omega_{\mathrm{m}})}}.$ (12)
The $w_{0}w_{a}\mathrm{CDM}$ model can be mapped in the same manner as
$q_{0}=-1+\frac{3}{2}\Omega_{\mathrm{m}}+\frac{3(1+w_{0})}{2}(1-\Omega_{\mathrm{m}}),$ (13)
and the corresponding $p_{\mathrm{age}}$ takes
$p_{\mathrm{age}}=\int_{0}^{1}\frac{\mathrm{d}a}{a\,\sqrt{\Omega_{\mathrm{m}}\,a^{-3}+(1-\Omega_{\mathrm{m}})a^{-3(1+w_{0}+w_{a})}e^{3w_{a}(1-a)}}}.$ (14)
Setting $w_{a}=0$ and $w_{0}=w$, the results are applicable to the $w\mathrm{CDM}$ model.
II.2 Cosmological datasets
In this section, we present an overview of various combinations of cosmological data used to construct the IDL, including SNe data from PantheonPlus and DESY5, BAO measurements from DESI DR2 and Sloan Digital Sky Survey (SDSS), and CC data.
II.2.1 SNe
SNe are known as the standard candles due to their uniform intrinsic luminosity. The SNe data records the apparent magnitude $m_{\mathrm{B}}(z)$. The luminosity distance $D_{\mathrm{L}}(z)$ can be derived from the magnitude–redshift relation
$m_{\mathrm{B}}(z)=5~\mathrm{lg}\frac{D_{\mathrm{L}}(z)}{10~\mathrm{pc}}+M_{\mathrm{B}}=5~\mathrm{lg}\frac{D_{\mathrm{L}}(z)H_{0}}{c}-5a_{\mathrm{B}},$ (15)
where $M_{\mathrm{B}}$ is the absolute magnitude of SNe. $a_{\mathrm{B}}\equiv-\frac{1}{5}[5~\mathrm{lg}(c/10~\mathrm{pc}/H_{0})+M_{\mathrm{B}}]$ is the intercept of the magnitude–redshift relation Eq. (15). It is evident that $H_{0}$ and $M_{\mathrm{B}}$ are degenerate in $a_{\mathrm{B}}$, making it impossible to constrain both $H_{0}$ and $M_{\mathrm{B}}$ with SNe data alone. In IDL, we incorporate BAO and CC (described below) to break this degeneracy, avoiding the dependence on external prior.
The chi-squared statistic for SNe is defined as follows
$\chi^{2}_{\mathrm{SNe}}=\Delta\mathbf{D}^{\mathrm{T}}\mathrm{C}^{-1}_{\mathrm{stat+sys}}\Delta\mathbf{D},$ (16)
where $\Delta{D}_{i}=\mu_{i}-\mu_{\mathrm{model}}(z_{i})$ is the distance moduli residual vector. Distance modulus is defined as $\mu\equiv m_{\mathrm{B}}-M_{\mathrm{B}}=5\mathrm{lg}(D_{\mathrm{L}}/10~\mathrm{pc})$. $\mathrm{C}^{-1}_{\mathrm{stat+sys}}$ denotes the inverse combined statistical and systematic covariance matrix of the SNe sample.
We use two different SNe datasets, including PantheonPlus [^3] PantheonPlus data available at https://github.com/PantheonPlusSH0ES/DataRelease and DESY5444DESY5 data available at https://github.com/des-science/DES-SN5YR.
The PantheonPlus SNe compilation consists of 1701 light curves of 1550 spectroscopically confirmed SNe Ia in the redshift range $0.001<z<2.26$ Scolnic et al. (2022). To reduce the effect of strong peculiar velocity dependence at low redshift Brout et al. (2022), only 1550 data points in the range $0.01<z<2.26$ are used.
The DESY5 SNe sample is the largest and deepest single sample survey to date consisting of 1635 SNe Ia Abbott et al. (2024); Vincenzi et al. (2024); Sánchez et al. (2024) distributed in $0.10<z<1.3$ complemented by 194 low-redshift SNe Ia spanning $0.025<z<0.1$ Hicken et al. (2009, 2012); Krisciunas et al. (2017); Foley et al. (2018). The DESY5 data provide the Hubble diagram redshift, corrected apparent magnitudes $m^{\mathrm{corr.}}_{\mathrm{B}}(z)$, distance moduli $\mu(z)$, and the corresponding distance moduli errors $\sigma_{ij}$ and related quantities. Since the distance moduli contain no information about $M_{\mathrm{B}}$, we follow Refs. Huang et al. (2025b); Camilleri et al. (2025) to use the corrected apparent magnitudes to construct the distance moduli residual as $\Delta\mathbf{D}=m_{\mathrm{B}}^{\mathrm{corr.}}-M_{\mathrm{B}}-5\mathrm{lg}({D_{\mathrm{L}}(z)}/{10~\mathrm{pc}})$ to constrain $M_{\mathrm{B}}$.
II.2.2 BAO
The BAO records the imprint of early-Universe plasma fluctuations in the late-time large-scale structure. BAO data provide three types of distances with respect to a fiducial cosmology, measured along different directions. Along the line-of-sight direction, BAO measures the Hubble distance,
$\frac{D_{\mathrm{H}}(z)}{r_{\mathrm{d}}}=\frac{c}{r_{\mathrm{d}}H(z)},$ (17)
and along the transverse direction, BAO measures the (comoving) angular diameter distance,
$\frac{D_{\mathrm{M}}(z)}{r_{\mathrm{d}}}=\frac{(1+z)D_{\mathrm{A}}(z)}{r_{\mathrm{d}}}=\frac{D_{\mathrm{L}}(z)}{(1+z)r_{\mathrm{d}}}.$ (18)
Taking the spherical average of both directions, BAO measures
$\frac{D_{\mathrm{V}}(z)}{r_{\mathrm{d}}}=\frac{[zD_{\mathrm{M}}(z)^{2}D_{\mathrm{H}}(z)]^{1/3}}{r_{\mathrm{d}}},$ (19)
where $r_{\mathrm{d}}$ is treated as a free parameter in our analysis to eliminate the model dependence introduced by the $r_{\mathrm{d}}$ prior in traditional IDL.
The chi-squared statistic for BAO takes
$\chi^{2}_{\mathrm{BAO}}=\mathbf{\Delta}^{\mathrm{T}}\mathrm{C}^{-1}_{\mathrm{BAO}}\mathbf{\Delta},$ (20)
where $\mathbf{\Delta}$ denotes the residual vector between the BAO data and the corresponding model values. $\mathrm{C}^{-1}_{\mathrm{BAO}}$ is the inverse covariance matrix of the BAO data.
Two BAO datasets from DESI [^5] DESI DR2 data available at https://github.com/CobayaSampler/bao_data/tree/master/desi_bao_dr2. and SDSS [^6] SDSS data available at https://svn.sdss.org/public/data/eboss/DR16cosmo/tags/v1_0_0/likelihoods/BAO-plus/ are used to delve into the role that BAO plays in the cosmological constraint, and these datasets are listed in Table 1.
We adopt the DESI DR2 BAO measurements which include $D_{\mathrm{M}}(z)/r_{\mathrm{d}},~D_{\mathrm{H}}(z)/r_{\mathrm{d}},~D_{\mathrm{V}}(z)/r_{\mathrm{d}}$ extracted from large scale structure traces across multiple distinct redshift bins Abdul Karim et al. (2025).
The SDSS BAO data are employed in this paper to cross-check our findings and connect with the existing literature. It includes two low-redshift data points 6dFGS Beutler et al. (2011), SDSS DR7 MGS Ross et al. (2015), and SDSS DR16 data Alam et al. (2021).
Table 1: The BAO samples utilized in this work including DESI DR2 data and SDSS data.
[…]
DESI DR2
BGS 0.30 $\cdots$ $\cdots$ 7.94 $\pm$ 0.08
LRG1 0.51 13.59 $\pm$ 0.17 21.86 $\pm$ 0.43 $\cdots$
LRG2 0.71 17.35 $\pm$ 0.18 19.46 $\pm$ 0.33 $\cdots$
LRG3 + ELG1 0.93 21.58 $\pm$ 0.15 17.64 $\pm$ 0.19 $\cdots$
ELG2 1.32 27.60 $\pm$ 0.32 14.18 $\pm$ 0.22 $\cdots$
QSO 1.48 30.51 $\pm$ 0.76 12.82 $\pm$ 0.52 $\cdots$
[…]
SDSS
6dFGS 0.106 $\cdots$ $\cdots$ 0.336 $\pm$ 0.015
MGS 0.15 $\cdots$ $\cdots$ 4.47 $\pm$ 0.17
LRG 0.70 17.86 $\pm$ 0.33 19.33 $\pm$ 0.53 $\cdots$
ELG 0.85 $\cdots$ $\cdots$ $18.33^{+0.57}_{-0.62}$
QS0 1.48 30.69 $\pm$ 0.80 13.26 $\pm$ 0.55 $\cdots$
[…]
Table 2: 32 CC data points with $1\sigma$ error.
z{z} $H(z)~\mathrm{[km~s^{-1}~Mpc^{-1}}]$} Ref.
0.09 69 $\pm$ 12 Jimenez et al. (2003)
0.17 83 $\pm$ 8 Simon et al. (2005)
0.27 77 $\pm$ 14 Simon et al. (2005)
0.4 95 $\pm$ 17 Simon et al. (2005)
0.9 117 $\pm$ 23 Simon et al. (2005)
1.3 168 $\pm$ 17 Simon et al. (2005)
1.43 177 $\pm$ 18 Simon et al. (2005)
1.53 140 $\pm$ 14 Simon et al. (2005)
1.75 202 $\pm$ 40 Simon et al. (2005)
0.48 97 $\pm$ 62 Stern et al. (2010)
0.88 90 $\pm$ 40 Stern et al. (2010)
0.1791 75 $\pm$ 4 Moresco et al. (2012)
0.1993 75 $\pm$ 5 Moresco et al. (2012)
0.3519 83 $\pm$ 14 Moresco et al. (2012)
0.5929 104 $\pm$ 13 Moresco et al. (2012)
0.6797 92 $\pm$ 8 Moresco et al. (2012)
0.7812 105 $\pm$ 12 Moresco et al. (2012)
0.8754 125 $\pm$ 17 Moresco et al. (2012)
1.037 154 $\pm$ 20 Moresco et al. (2012)
0.07 69 $\pm$ 19.6 Zhang et al. (2014b)
0.12 68.6 $\pm$ 26.2 Zhang et al. (2014b)
0.2 72.9 $\pm$ 29.6 Zhang et al. (2014b)
0.28 88.8 $\pm$ 36.6 Zhang et al. (2014b)
1.363 160 $\pm$ 36.6 Moresco (2015)
1.965 186.5 $\pm$ 50.4 Moresco (2015)
0.3802 83 $\pm$ 13.5 Moresco et al. (2016)
0.4004 77 $\pm$ 10.2 Moresco et al. (2016)
0.4247 87.1 $\pm$ 11.2 Moresco et al. (2016)
0.4497 92.8 $\pm$ 12.9 Moresco et al. (2016)
0.4783 80.9 $\pm$ 9 Moresco et al. (2016)
0.47 89 $\pm$ 49.6 Ratsimbazafy et al. (2017)
0.80 113.1 $\pm$ 28.5 Jiao et al. (2023)
II.2.3 CC
CC measurements are incorporated to break the $r_{\mathrm{d}}$ and $H_{0}$ degeneracy in BAO data, serving a role analogous to the $r_{\mathrm{d}}$ prior in traditional IDL. It directly measures the Hubble parameter through massive passively evolving galaxies with old stellar populations and very low star formation rates Moresco (2023). By comparing two galaxies formed at the same time but separated by a small redshift interval, the Hubble parameter $H(z)$ can be calculated using the differential age method Jimenez and Loeb (2002), without assuming any cosmological model,
$H(z)=-\frac{1}{1+z}\frac{\mathrm{d}z}{\mathrm{d}t},$ (21)
where ${\mathrm{d}z}/{\mathrm{d}t}$ is the derivative of redshift with respect to the look-back time. Its nature as a direct measurement of $H_{0}$ makes it highly sensitive to cosmological parameters.
As CC data are not entirely independent, we incorporate correlations through the use of a covariance matrix when available. We adopt 32 CC data points, listed in Table 2, including 15 correlated measurements with the corresponding covariance matrix compiled in Ref. Moresco et al. (2020) and 17 uncorrelated ones. For the correlated measurements, we adopt the covariance matrix as described in Refs. Moresco et al. (2020, 2022),
$\mathrm{C}_{ij}=\mathrm{C}^{\mathrm{stat}}_{ij}+\mathrm{C}^{\mathrm{sys}}_{ij},$ (22)
where $\mathrm{C}^{\mathrm{stat}}_{ij}$ represents the diagonal covariance matrix containing statistical uncertainties, $\mathrm{C}^{\mathrm{sys}}_{ij}$ accounts for systematic effects, including those from stellar population models, metallicity assumptions, and potential contamination from residual young stellar components.
In this work, we adopt the recommended covariance matrix combination from Moresco (2021),
$\mathrm{C}^{\mathrm{tot}}_{ij}=\mathrm{C}^{\mathrm{stat}}_{ij}+\mathrm{C}_{ij}^{\mathrm{met}}+\mathrm{C}_{ij}^{\mathrm{sps,ooo}}+\mathrm{C}_{ij}^{\mathrm{IMF}},$ (23)
where $\mathrm{C}_{ij}^{\mathrm{met}}$ denotes the contribution from uncertainties in stellar metallicity estimates. $\mathrm{C}_{ij}^{\mathrm{sps,ooo}}$ accounts for uncertainties arising from the odd one out stellar population synthesis (SPS) model. $\mathrm{C}_{ij}^{\mathrm{IMF}}$ reflects uncertainties due to the adopted initial mass function (IMF). For a detailed discussion of CC data, we refer the reader to Ref. Moresco et al. (2020).
For data points not discussed in Ref. Moresco et al. (2020), only the diagonal components are considered. The final $\chi^{2}$ is
$\chi_{\mathrm{CC}}^{2}=\sum_{i=1}^{17}\left[\frac{H_{\mathrm{obs}}\left(z_{i}\right)-H_{\mathrm{model}}\left(z_{i}\right)}{\sigma_{H,_{i}}}\right]^{2}+\Delta\mathbf{H}^{\rm T}\mathrm{C}_{\mathrm{tot}}^{-1}\Delta\mathbf{H},$ (24)
where $\sigma_{\mathrm{H}}$ are the diagonal components mentioned above, and $\Delta\mathbf{H}$ is the residual vector between the CC data and the model values.
We consider four datasets, namely DESY5 + DESI DR2 + CC, DESY5 + SDSS + CC, PantheonPlus + DESI DR2 + CC, and PantheonPlus + SDSS + CC. The final combined likelihood, $2\mathrm{ln}\mathcal{L}=\chi^{2}_{\mathrm{SNe}}+\chi^{2}_{\mathrm{BAO}}+\chi^{2}_{\mathrm{CC}}$, is used to constrain the PAge model with five free parameters $\{\eta,~p_{\mathrm{age}},~M_{\mathrm{B}},~H_{0},~r_{\mathrm{d}}\}$. Then we employ the Markov Chain Monte Carlo (MCMC) sampler emcee [^7] https://github.com/dfm/emcee Foreman-Mackey et al. (2013) to explore the parameter space and obtain best-fit parameters through $\chi^{2}$ minimization. For comparison, we also consider the $\Lambda\mathrm{CDM}$ model, with four free parameters $\{\Omega_{\mathrm{m}},~M_{\mathrm{B}},~H_{0},~r_{\mathrm{d}}\}$. Uniform priors are adopted for the following parameter ranges $\eta\in[-2,2],~p_{\mathrm{age}}\in[0.15,2],~M_{\mathrm{B}}\in[-21,-19],~H_{0}\in[60,80]~\mathrm{km~s^{-1}~Mpc^{-1}},~r_{\mathrm{d}}\in[120,180]~\mathrm{Mpc},~\Omega_{\mathrm{m}}\in[0.2,0.4]$. Posterior samples are analyzed and visualized using GetDist [^8] https://github.com/cmbant/getdist Lewis (2025). For model comparison, we use the MCEvidence [^9] https://github.com/yabebalFantaye/MCEvidence algorithm as described in Ref. Heavens et al. (2017) to compute the Bayesian evidence directly from the MCMC chain.
Table 3: Cosmological constraints on the PAge and $\Lambda\mathrm{CDM}$ models based on the DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC datasets. Here, $H_{0}$ and $r_{\mathrm{d}}$ are in units of $\mathrm{km~s^{-1}~Mpc^{-1}}$ and $\mathrm{Mpc}$.
Dataset $\eta$ $p_{\mathrm{age}}$ $\Omega_{\mathrm{m}}$ $M_{\mathrm{B}}$ $H_{0}$ $r_{\mathrm{d}}$
PAge
DESY5+DESI DR2+CC 0.23 $\pm$ 0.04 0.94 $\pm$ 0.01 $\cdots$ $19.36\pm$ 0.07 67.91 $\pm$ 2.33 $145.0^{+4.5}_{-5.2}$
DESY5+SDSS+$C$ 0.16 $\pm$ 0.06 0.95 $\pm$ 0.01 $\cdots$ $19.35\pm$ 0.07 67.74 $\pm$ 2.36 $143.5^{+4.6}_{-5.3}$
PantheonPlus+DESI DR2+CC 0.31 $\pm$ 0.06 0.96 $\pm$ 0.01 $\cdots$ $19.38\pm$ 0.07 68.97 $\pm$ 2.42 $145.0^{+4.5}_{-5.2}$
PantheonPlus+SDSS+$C$ 0.25 $\pm$ 0.07 0.96 $\pm$ 0.02 $\cdots$ $19.39\pm$ 0.08 68.58 $\pm$ 2.49 $143.4^{+4.6}_{-5.3}$
$\bm{\Lambda\mathrm{CDM}}$
DESY5+DESI DR2+CC $\cdots$ $\cdots$ 0.32 $\pm$ 0.01 $19.35\pm$ 0.07 68.99 $\pm$ 2.35 $145.2^{+4.5}_{-5.1}$
DESY5+SDSS+$C$ $\cdots$ $\cdots$ 0.35 $\pm$ 0.01 $19.38\pm$ 0.07 67.46 $\pm$ 2.37 $143.4^{+4.6}_{-5.3}$
PantheonPlus+DESI DR2+CC $\cdots$ $\cdots$ 0.30 $\pm$ 0.01 $19.37\pm$ 0.07 69.75 $\pm$ 2.39 $145.0^{+4.5}_{-5.2}$
PantheonPlus+SDSS+$C$ $\cdots$ $\cdots$ 0.32 $\pm$ 0.02 $19.39\pm$ 0.08 69.08 $\pm$ 2.51 $143.2^{+4.6}_{-5.3}$
III RESULTS AND DISCUSSION
Our primary goal is to derive a model-independent estimate of $H_{0}$, as well as other cosmological parameters $M_{\mathrm{B}}$ and $r_{\mathrm{d}}$, using the PAge-improved IDL. The summary of the full cosmological results is presented in Table 3, along with the corresponding $\Lambda\mathrm{CDM}$ result for comparison. The one- and two-dimensional posterior distributions of the cosmological parameters in the PAge model are presented in Fig. 2. For the DESY5+DESI DR2+CC datasets, the PAge model yields
$H_{0}=67.91\pm 2.33~\mathrm{km~s^{-1}~Mpc^{-1}},$ (25)
at the $1\sigma$ confidence level. Our estimate agrees with the Planck 2018 result of $H_{0}=67.4\pm 0.5~\mathrm{km~s^{-1}~Mpc^{-1}}$ Aghanim et al. (2020). However, a mild tension persists at the $2.0\sigma$ level in comparison with the local SH0ES measurement of $H_{0}=73.04\pm 1.04~\mathrm{km~s^{-1}~Mpc^{-1}}$ Riess et al. (2022). Other datasets exhibit the same agreement with the Planck result and a similar tension level with the SH0ES measurements. Specifically, DESY5+SDSS+CC datasets show a $2.1\sigma$ discrepancy, while PantheonPlus+DESI DR2+CC and PantheonPlus+SDSS+CC datasets exhibit $1.6\sigma$ and $1.7\sigma$ discrepancies, respectively. Our results show larger uncertainty than other traditional IDLs Camilleri et al. (2025); Aubourg et al. (2015); Macaulay et al. (2019) due to the substitution of the Gaussian $r_{\mathrm{d}}$ prior with the more uncertain CC measurements.
From DESY5+DESI DR2+CC, we obtain $M_{\mathrm{B}}=-19.36\pm 0.07$, consistent with the Gaussian processes estimate $M_{\mathrm{B}}=-19.31^{+0.09}_{-0.11}$ Favale et al. (2023), with $M_{\mathrm{B}}=-19.303$ derived from tip of the red giant branch with full sample in Hubble Space Telescope or James Webb Space Telescope Li et al. (2025f), and $1.43\sigma$ differing from the SH0ES result $M_{\mathrm{B}}=-19.253\pm 0.027$ Riess et al. (2022). For sound horizon, we find $r_{\mathrm{d}}=145.0^{+4.5}_{-5.2}~\mathrm{Mpc}$, in agreement with Planck result $r_{\mathrm{d}}=147.05\pm 0.30~\rm Mpc$ and $r_{\mathrm{d}}=146.0^{+4.2}_{-5.1},~141.9^{+5.6}_{-4.9}~\mathrm{Mpc}$ reported in Refs. Gómez-Valent (2022); Favale et al. (2023). As shown in Fig. 2, despite the constraints on $M_{\mathrm{B}}$ and $r_{\mathrm{d}}$ being broad and remain $\sim 1\sigma$ consistent with Planck and SH0ES, the value of $H_{0}$ differs from SH0ES by $2.0\sigma$. This higher discrepancy stems from $a_{\mathrm{B}}$ offset between DESY5 and PantheonPlus and difference in $H_{0}r_{\mathrm{d}}$ inferred from DESI and SDSS. As reported in Ref. Efstathiou (2025), DESY5 and PantheonPlus exhibit an offset of $\sim 0.04$ magnitude in their overlap. Moreover, Ref. Huang et al. (2025b) finds that $5a_{\mathrm{B}}$ for DESY5 is higher than that for PantheonPlus. From Eq. (15), at fixed $M_{\mathrm{B}}$, a larger value of $5a_{\mathrm{B}}$ implies smaller $H_{0}$, thereby increasing the tension with SH0ES. As shown in the $H_{0}-M_{\mathrm{B}}$ plane of Fig. 2, holding $M_{\mathrm{B}}$ fixed, DESY5 favors a lower value of $H_{0}$, consistent with Refs. Colgáin and Sheikh-Jabbari (2024); Notari et al. (2024); Adame et al. (2025). Note that the value of $H_{0}r_{\mathrm{d}}$ inferred from DESI is higher than those from SDSS Mukherjee and Sen (2024); Adame et al. (2025), which would, for fixed $r_{\mathrm{d}}$, raise the inferred $H_{0}$. Nevertheless, this effect is smaller than the effect of SNe as shown in $H_{0}-M_{\mathrm{B}}$ and $H_{0}-r_{\mathrm{d}}$ planes across four datasets in Fig. 2. Consequently, although $M_{\mathrm{B}}$ is within $1.4\sigma$ of SH0ES, the inferred $H_{0}$ differs from the SH0ES value by $2.0\sigma$.
Figure 2: Constraints on the PAge model from DESY5+DESI DR2+CC (green), DESY5+SDSS+$C$ (blue, dotted), PantheonPlus+DESI DR2+CC (purple), and PantheonPlus+SDSS+CC (brown, dotted). Planck 2018 (pink) and SH0ES (grey) benchmarks are included to illustrate the $H_{0}$ tension both from the direct $H_{0}$ constraints and from those inferred via $M_{\mathrm{B}}$ and $r_{\mathrm{d}}$. The SH0ES-equivalent $r_{\mathrm{d}}$ and Planck-equivalent $M_{\mathrm{B}}$ are obtained by equating $H_{0}r_{\mathrm{d}}$ and $a_{\mathrm{B}}$, respectively. Four datasets are presented in the $M_{\mathrm{B}}$$H_{0}$ plane to illustrate the shift in the intercept of the magnitude-redshift relation vividly.
Figure 3: The constraints on $\eta$ and $p_{\mathrm{age}}$ in the PAge model obtained from DESY5+DESI DR2+CC (green), DESY5+SDSS+$C$ (blue), PantheonPlus+DESI DR2+CC (purple), and PantheonPlus+SDSS+CC (pink). The DESI DR2 and Planck 2018 results are introduced for comparison. Since $\Lambda\mathrm{CDM}$ lacks one parameter compared to the PAge model, it is mapped as a straight line in the parameter space.
Figure 4: The BAO distance for the PAge (blue), $w_{0}w_{a}\mathrm{CDM}$ (green), and $w\mathrm{CDM}$ (orange) models in DESY5 + DESI DR2 + CC. These distances are normalized by the best-fit $\Lambda\mathrm{CDM}$ model (black) in DESY5+DESI+$C$. The data of DESI (purple, filled) are drawn with error bars.
Figure 5: The reconstructed BAO distance for the PAge (green), MAPAge (red), incorporating DESI DR2 data. The shaded bands denote the corresponding $1\sigma$ uncertainties.
Figure 6: Bayesian model comparison of PAge with $\Lambda\rm CDM$ (left panel) and $w_{0}w_{a}\mathrm{CDM}$ model (right panel). Positive values indicate a preference for the PAge model in each dataset. The impact of prior widths for $\eta$ and $p_{\mathrm{age}}$ is assessed by broadening the original prior width $d$ to $d^{*}=\alpha d$, with $\alpha$ denoting the broadening factor.
After examining how the choice of SNe and BAO datasets influences the cosmological parameters, we now focus on the approximation parameters of the PAge model itself, namely $\eta$ and $p_{\mathrm{age}}$. These two parameters play a central role in offering insights into potential late-time modifications to cosmic expansion. The posterior distributions in the $(p_{\mathrm{age}},\eta)$ plane for the four datasets are shown in Fig. 3. For comparison, previous results from $\Lambda\mathrm{CDM}$ Aghanim et al. (2020) and DESI DR2 Abdul Karim et al. (2025) are mapped into the PAge model. The mapping procedure is detailed in Eqs. (6), (11), (12), (13), and (14). Note that the PAge model contains one additional parameter compared to $\Lambda\mathrm{CDM}$, which causes the mapped region of $\Lambda\mathrm{CDM}$ in the $(p_{\mathrm{age}},\eta)$ space to appear as a line. We adopt the DESI DR2 results for $w\mathrm{CDM}$ and $w_{0}w_{a}\mathrm{CDM}$ models Abdul Karim et al. (2025). As shown in Fig. 3, it is evident that our results deviate from the $\Lambda\mathrm{CDM}$ model, showing a lower value of $\eta$ for the DESY5 dataset. In contrast, for the PantheonPlus samples, different BAO combinations remain consistent with $\Lambda\mathrm{CDM}$.
To better understand these results, we reconstruct the BAO distance predicted by the PAge, $w\mathrm{CDM}$ and $w_{0}w_{a}\mathrm{CDM}$ in Fig. 4, based on our best-fit $\Lambda\mathrm{CDM}$ model in the DESY5+DESI DR2+CC datasets. The PAge model exhibits a distinctive trend driven by the DESI DR2 data. In comparison to the $\Lambda\mathrm{CDM}$ model, it predicts slightly enhanced BAO distances at very low redshifts, suppressed distances at intermediate redshifts, and elevated values again at high redshifts across all configurations. While the $w\mathrm{CDM}$ model attempts to capture this behavior, it fails to reproduce the variation across the relevant redshift range. In contrast, the $w_{0}w_{a}\mathrm{CDM}$ model, which incorporates dynamically evolving equation of state, is more capable of reproducing this characteristic trend.
These findings suggest that the DESI DR2 data may not be fully compatible with the $\Lambda\mathrm{CDM}$ model, which lack the flexibility to capture such rapid variations in the expansion history, thereby hinting at a preference for dynamical dark energy. PAge provides a model-independent description of the underlying behaviors of the late-time Universe. However, it should be noted that the PAge model rests on specific assumptions. Thus, PAge may not capture inhomogeneous cosmologies, nor scenarios that violate the assumptions (i) and (ii).
Then, we fit the MAPAge model to DESY5+DESI DR2+CC instead of PAge, to assess the impact of the chosen truncated order in Eq. (1). Our final constraints are: $\eta_{1}=-0.01\pm 0.09,~\eta_{2}=0.65\pm 0.22,~p_{\mathrm{age}}=0.93\pm 0.01,~H_{0}=67.57\pm 2.33~\mathrm{km~s^{-1}~Mpc^{-1}},~r_{\mathrm{d}}=144.7^{+4.5}_{-5.2}~\mathrm{Mpc}$. The constrains of $p_{\mathrm{age}},~H_{0},~M_{\mathrm{B}},~r_{\mathrm{d}}$ agree with PAge, except for $\eta~(\eta_{1})$, which shifts toward more negative values and has a larger uncertainty in MAPAge, caused by the added cubic term. Furthermore, the reconstructed BAO distances $D_{\mathrm{H}}/r_{\mathrm{d}},~D_{\mathrm{M}}/r_{\mathrm{d}},~D_{\mathrm{v}}/r_{\mathrm{d}}$ are shown in Fig. 5. The results exhibit no significant difference from PAge in either central values or uncertainties. The enlarged uncertainty in $\eta~(\eta_{1})$ minimally affects $D_{\mathrm{H}}/r_{\mathrm{d}},~D_{\mathrm{M}}/r_{\mathrm{d}},~D_{\mathrm{v}}/r_{\mathrm{d}}$, demonstrating stability with respect to the chosen truncated order in Eq. (1).
To quantitatively evaluate the performance of the PAge model relative to the $\Lambda\mathrm{CDM}$ model, we perform Bayesian inference across different datasets. The MAPAge does not show significant different with PAge, as demonstrated by Bayesian inference. Therefore, for simplicity , we just focus on PAge model. Specifically, the Bayes factors $\mathrm{ln}\mathcal{B}_{ij}$ are calculated in logarithmic space,
$\mathrm{ln}\mathcal{B}_{ij}=\mathrm{ln}Z_{i}-\mathrm{ln}Z_{j},$ (26)
where $Z_{i}$ and $Z_{j}$ are Bayesian evidence of PAge and $\Lambda\mathrm{CDM}$ models, respectively. The Bayesian evidence is defined as
$Z\equiv\int P(D\mid\theta,M)\,P(\theta\mid M)\,P(M)\,\mathrm{d}\theta,$ (27)
where $P(D\mid\theta,M)$ is the likelihood of the data $D$ given the parameters $\theta$ and model $M$, $P(\theta\mid M)$ is the prior probability of $\theta$ under $M$, and $P(M)$ is the prior of $M$. The extent of model preference is assessed using the Jeffreys scale, as described in Ref. Trotta (2008).
  • $|\mathrm{ln}\mathcal{B}_{ij}|<1$, inconclusive evidence.
  • $1\leq|\mathrm{ln}\mathcal{B}_{ij}|<2.5$, weak evidence.
  • $2.5\leq|\mathrm{ln}\mathcal{B}_{ij}|<5$, moderate evidence.
  • $5\leq|\mathrm{ln}\mathcal{B}_{ij}|<10$, strong evidence.
  • $10\leq|\mathrm{ln}\mathcal{B}_{ij}|$, decisive evidence.
Given the impact of prior width on our results, we follow Refs. Patel et al. (2024); Sola Peracaula et al. (2019) and broaden the original prior width $d$ by a factor $\alpha$, so that the new prior width is $d^{*}=\alpha d$. In the case of uniform priors Patel et al. (2024); Sola Peracaula et al. (2019), the new Bayes factors $\ln\mathcal{B}^{*}_{ij}$ are modified into
$\ln\mathcal{B}_{ij}^{*}=\ln\left[\frac{(\alpha_{\nu_{1}}\alpha_{\nu_{2}}\alpha_{\nu_{3}}\dots)_{j}}{(\alpha_{\nu_{1}}\alpha_{\nu_{2}}\alpha_{\nu_{3}}\dots)_{i}}\right]+\ln\mathcal{B}_{ij},$ (28)
where $(\alpha_{\nu_{1}}\alpha_{\nu_{2}}\alpha_{\nu_{3}}\dots)_{i}$ denotes the product of prior-width broadening factors $\alpha_{\nu_{k}}$ for parameters $\nu_{k}$ in PAge and $(\alpha_{\nu_{1}}\alpha_{\nu_{2}}\alpha_{\nu_{3}}\dots)_{j}$ denotes the corresponding product for $\Lambda\mathrm{CDM}$. We consider only the impact of the prior widths of $\eta,~p_{\mathrm{age}}$ in PAge. For simplicity, we set $\alpha_{\eta}=~\alpha_{p_{\mathrm{age}}}=\alpha$. So the Eq. (28) becomes
$\ln\mathcal{B}_{ij}^{*}=\ln\mathcal{B}_{ij}-2\ln\alpha.$ (29)
In the left panel of Fig. 6, we present the final model comparison results, with the PAge model assessed relative to the fiducial $\Lambda\mathrm{CDM}$ across four datasets. Positive values indicate a preference for the PAge model, while negative values indicate a preference for the $\Lambda\mathrm{CDM}$ model. Specifically, when $\alpha=1$ we find $\mathrm{ln}\mathcal{B}_{ij}=9.17,~4.81,~1.01,~\text{and}~2.71$ for DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively. It is clear that the PAge model consistently provides a better fit than the standard $\Lambda\mathrm{CDM}$ across all datasets. In particular, for the DESY5+DESI DR2+CC datasets, the Bayes factor is 9.17, indicating nearly decisive evidence in favor of the PAge model over the $\Lambda\mathrm{CDM}$ model. Varying $\alpha$ can switch the preferred model. When $\alpha\geq 12,~2,~4$ for DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively, the Bayes factor favors $\Lambda\mathrm{CDM}$. By contrast, for DESY5+DESI DR2+CC, the preference flips for $\alpha\geq 98$.
Moreover, to assess the performance of PAge compared to $w_{0}w_{a}\mathrm{CDM}$, we perform Bayesian inference in the same manner. The prior ranges used for $w_{0}w_{a}\mathrm{CDM}$ in this work are $w_{0}\in[-3,1],~w_{a}\in[-3,2],~M_{\mathrm{B}}\in[-21,-19],~H_{0}\in[60,80]~\mathrm{km~s^{-1}~Mpc^{-1}},~r_{\mathrm{d}}\in[120,180]~\mathrm{Mpc},~\Omega_{\mathrm{m}}\in[0.2,0.4]$. We also explore the impact of prior width by broadening the prior widths on $\eta$ and $p_{\mathrm{age}}$. The Bayes factors take the same form as Eq. (29), where $i$ denotes PAge and $j$ denotes $w_{0}w_{a}\mathrm{CDM}$. The final results are shown in the right panel of Fig. 6. For $\alpha=1$, we find $\ln\mathcal{B}_{ij}=-1.25,~0.32,~1.31,~\text{and}~2.41$ for DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively. These results indicate at most weak evidence in favor of $w_{0}w_{a}\mathrm{CDM}$ model over PAge model. For DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, the evidence shifts in favor of $w_{0}w_{a}\mathrm{CDM}$ when $\alpha\geq 2,~2,~4$, respectively. By contrast, for DESY5+DESI DR2+CC, PAge is favored only when $\alpha\leq 0.5$. These conclusions also depend on the chosen prior width for $w_{0},~w_{a}$, which are held fixed here for simplicity. Widening those prior widths on $w_{0}$ and $w_{a}$ tends to strengthen the evidence in favor of the PAge model.
IV CONCLUSION
The tension between the early-Universe and late-Universe measurements of $H_{0}$ has become one of the most significant challenges in modern cosmology. Recently, DESI DR1 Adame et al. (2025) and DR2 Abdul Karim et al. (2025) have further sharpened this tension with addition hints of dynamical dark energy in $3.9\sigma$ and $4.2\sigma$ level. In the context of the $H_{0}$ tension, numerous theoretical models have been proposed, although many models face corresponding no-go arguments. Similarly, the hint of dynamical dark energy has been critically examined through the reanalyzes of the DESY5 SNe data Huang et al. (2025b); Efstathiou (2025); Vincenzi et al. (2025); Gialamas et al. (2025b) and the DESI BAO measurements Zheng et al. (2025); Mukherjee and Sen (2025); Liu et al. (2024).
To provide a model-independent perspective, we employ the PAge-improved IDL in this work to constrain cosmological parameters. Our work is based on four datasets, namely DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, allowing us to investigate the respective roles of SNe and BAO data in constraining cosmological parameters. For DESY5+DESI DR2+CC, we find $H_{0}=67.91\pm 2.33~\mathrm{km~s^{-1}~Mpc^{-1}}$, which agrees with the Planck 2018 result, but exhibits a mild $2.0\sigma$ tension with the SH0ES measurement. The other datasets also remain consistent with Planck 2018. Compared to SH0ES, the resulting tensions are $2.1\sigma$, $1.6\sigma,$ and $1.7\sigma$ for DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively. In our results, the DESY5 data tend to favor a lower value of $H_{0}$ and a higher value of $M_{\mathrm{B}}$, which correspond to a shift in the intercept of the magnitude–redshift relation within the $H_{0}$$M_{\mathrm{B}}$ parameter space. Additionally, the DESI DR2 data indicate a preference for a higher value of $r_{\mathrm{d}}$ compared to SDSS data, with a difference of approximately 2 Mpc.
Mapping specific cosmological models into the PAge parameter space $(p_{\mathrm{age}},\eta)$, our results align with the DESI DR2 hints of dynamical dark energy and deviate from $\Lambda\mathrm{CDM}$. In particular, the reconstructed BAO distance of the PAge model reveals a special trend compared to best-fit $\Lambda\mathrm{CDM}$. It shows a slight enhancement in very low redshift, suppression at intermediate redshift, and an upturn at high redshift, which may suggest that the $\Lambda\mathrm{CDM}$ model is insufficient to fully reconcile the observed data. Furthermore, the Bayesian model comparison provides evidence in favor of the PAge model over $\Lambda\mathrm{CDM}$, with Bayes factors of 9.17, 4.81, 1.01, and 2.71 in DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, PantheonPlus+SDSS+CC respectively. It gives nearly decisive evidence in favor of the PAge model over $\Lambda\mathrm{CDM}$ in the DESY5+DESI DR2+CC combination. Comparing PAge with $w_{0}w_{a}\mathrm{CDM}$, the Bayes factors are 1.25,0.32,1.31,-1.25,~0.32,~1.31, and $2.41$ for DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively, showing at most weak evidence in favor of $w_{0}w_{a}\mathrm{CDM}$. The interpretation of these comparisons is influenced by the adopted prior width. The wider prior width for PAge parameters, the weaker the evidence in favor of PAge.
In future work, we aim to explore the data in more detail by performing a redshift-binned analysis. Motivated by the connection between the expansion history and the dark-energy equation of state, we will further investigate the occurrence of unphysical phantom behaviors Caldwell et al. (2003); Vikman (2005) using the PAge model. It should be noted that the PAge model is essentially a Taylor expansion, which may suffer from singularities as discussed in Ref. Zhang et al. (2017b) when reconstructing the equation-of-state parameter. We are going to calibrate the Taylor expansion with alternatives beyond the early-matter-dominated Universe used in the PAge model, to avoid the singularities. These efforts may provide new insights into the nature of dark energy and further clarify the physical origin of the $H_{0}$ tension.
Acknowledgements.
We thank Sheng-Han Zhou and Yi-Min Zhang for their helpful discussions. This work was supported by the National SKA Program of China (Grants No. 2022SKA0110200 and No. 2022SKA0110203), the National Natural Science Foundation of China (Grants No. 12533001, No. 12575049, No. 12473001, No. 11975072, No. 11875102, No. 11835009, No. 12422502, No. 12105344, No. 12235019, No. 12447101, No. 12073088, No. 11821505, No. 11991052, and No. 11947302), the China Manned Space Program (Grants No. CMS-CSST-2025-A01 and No. CMS-CSST-2025-A02), the National 111 Project (Grant No. B16009), and the National Key Research and Development Program of China (Grants No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501).
DATA AVAILABILITY
The data that support the finding of this article are openly available Ling et al. .